OpenStudy (calculusxy):

A store held a going-out-of business sale where everything was 35% off. Luisa spent a total of $125.45. What would she have spent on the same items if there had been no sale?

OpenStudy (calculusxy):

@perl @mathstudent55

OpenStudy (calculusxy):

0.65x=125.45 ?

OpenStudy (perl):

Let x = undiscounted original price We can use the equation x ( 1 - 35/100) = 125.45

OpenStudy (perl):

correct

OpenStudy (perl):

solve for x

OpenStudy (calculusxy):

x=193

OpenStudy (calculusxy):

I have another question, that I just want to make sure that I have the correct answer. @perl Can you check?

OpenStudy (calculusxy):

John is buying a gold chain that had an original price of $950 but is advertised as 45% off. Sales tax of 9.50% applies to the discounted price. How much will John pay for the chain?

OpenStudy (calculusxy):

I got the answer of $572. 14.

OpenStudy (perl):

we have the same equation, but this time you multiply twice

OpenStudy (perl):

original price = 950 discounted price = 950 * ( 1 - .45) discounted price & sales tax = 950(1-.45) ( 1 + .095)

OpenStudy (perl):

which is equal to 572.1375 ~ 572.14 I get the same answer as you

OpenStudy (calculusxy):

Ok thank you. Peace :)